OpenStudy (anonymous):

Find the values of the six trigonometric functions of θ. Function Value: Sin θ = 0 and the constraint is: π/2≤θ≤3π/2

hartnn (hartnn):

in the range π/2≤θ≤3π/2, sin θ=0 when θ=π now cos π=-1 tan π=0 cot π=1/tan π=infinity sec π=1/cos π=-1 cosec π=1/ sin π= infinity